Rogers age is 1 1/3 times Stevens age. Eight years ago, Rogers age was twice Stevens age. Find Rogers age.
Answer by stanbon(48558)

Rogers age is 1 1/3 times Stevens age. Eight years ago, Rogers age was twice Stevens age. Find Rogers age.
------------------
Ages NOW:
Let Steven's age be "x".
Roger is now (4/3)x yrs old.
--------------------
Eight years AGO:
Steven was "x-8" years old.
Roger was (4/3)x-8 years old.
--------------------
EQUATION:
(4/3)x-8=2(x-8)
(4/3)x-8=2x-16
(4/3)x-(6/3)x=8-16
(-2/3) x =-8
x=12
Roger's age NOW is (4/3)(12)=16 years
Cheers,
Stan H.